University Theater sold 494 tickets for a play. Tickets cost $25 per adult and $10 per senior citizen. If total reciepts were $7100, how many senior citizen tickets were sold?
Answer by rfer(16322)

a+s=494 a=494-s
25a+10s=7100
25(494-s)+10s=7100
12350-25s+10s=7100
-15s=-5250
s=350
350*10=$3500
494-350=144 adults
144*25=$3600